OpenStudy (anonymous):

if sin o= 1/2....what are the exact ratios for the other five trig ratios?

jimthompson5910 (jim_thompson5910):

remember that sin^2 + cos^2 = 1 so sin^2 + cos^2 = 1 (1/2)^2 + cos^2 = 1 1/4 + cos^2 = 1 cos^2 = 1 - 1/4 cos^2 = 3/4 cos = sqrt(3/4) ... see note below cos = sqrt(3)/2 Note: the sign of the square root will depend on where the angle theta is. If theta is in quadrants 1 or 4, then cosine is positive here. If theta is in quadrants 2 or 3, then cosine is negative here

jimthompson5910 (jim_thompson5910):

once you know sine and cosine, you can find tangent since tan = sin/cos

jimthompson5910 (jim_thompson5910):

then you would use these 3 to find csc, sec, and cot

OpenStudy (luigi0210):